OpenStudy (anonymous):

infauzan help! There is an outbreak of the flu in a small town. The number of citizens who have been exposed to the new flu virus is given by the function F(t) = 11,000(1 − e−0.8t ) where t is the time in days after the initial outbreak. How many citizens have been exposed three days after the outbreak?

OpenStudy (anonymous):

Plugging this into a calculator results in 10002. \[F(t) = 11000(1-e^{-0.8(t)})\] \[F(3) = 11000(1-e ^{-.8(3)})=10002\]

OpenStudy (anonymous):

it's a good answer, i think.